if x-17=25 then x= what?
Found 2 solutions by mananth, ewatrrr:
Answer by mananth(16949)

x-17=25
add 17 to both sides
x-17+17 = 25+17
x=42

Answer by ewatrrr(24785)

 
Hi
x-17=25
x-17 +17 = 25+17 |Adding 17 to both sides of the = sign, to isolate x
x = 42